# An additive two-level parallel variant of the DMRG algorithm with coarse-space correction

### v1

@@ 2.2. Alternating Minimization on the Tensor Train Parameter Space
$\hskip 41.25641pt\mathbb{X}_{\rm LR}^{i,(n+1)}=(\mathbb{X}^{(n+1)}_{1},\ldots,\mathbb{X}^{(n+1)}_{i-1},\widetilde{\mathbb{X}}_{i}^{(n)},\mathbb{X}^{(n)}_{i+1},\ldots,\mathbb{X}^{(n)}_{d})$
$\hskip 24.18501pt\hookrightarrow\mathbb{X}_{\rm LR}^{i+1,(n+1)}=(\mathbb{X}^{(n+1)}_{1},\ldots,\mathbb{X}^{(n+1)}_{i-1},\mathbb{X}^{(n+1)}_{i},\widetilde{\mathbb{X}}^{(n)}_{i+1},\ldots,\mathbb{X}^{(n)}_{d}),$
$\hskip 41.25641pt\text{where }\mathbb{X}^{(n+1)}_{i}=\mathbb{U}_{i},~{}\widetilde{\mathbb{X}}^{(n+1)}_{i-1}=\mathbb{S}_{i}\mathbb{V}_{i}.$
@@
Remark 12 (Well-Posedness of Successive DMRG Micro-iterations).
Notice that in order for the DMRG Algorithm 1 to be well-posed, we require that for a given initialization $\mathbb{X}_{\rm LR}^{0,(0)}\in\mathcal{V}_{\mathbb{r}}$, each subsequent iterate $\mathbb{X}_{\rm LR}^{j,(n)}$, $\mathbb{X}_{\rm RL}^{k,(n)}$  $j\in\{1,\ldots,d-1\},~{}k\in\{2,\ldots,d\}$, and $n\in\mathbb{N}$ should also be an element of $\mathcal{V}_{\mathbb{r}}$. For the specific cases of quadratic or Rayleigh quotient-based energy functionals (and in the two-site DMRG case, for a judicious choice of rank truncation parameter $\widetilde{\mathbb{r}}$), it can be shown that this indeed holds [17, 42] provided that the initialization is sufficiently close to the true, sought-after minimizer. For a convergence analysis of the one-site DMRG algorithm, we refer interested readers to [36, 42].
@@ 3. Additive Two-Level DMRG with Coarse Space Correction
Equipped with the tools and notions presented in Section 2, we are now ready to present our proposal for the additive two-level DMRG algorithm with coarse space correction, or A2DMRG for short. Prior to presenting the details, let us first present a high-level intuitive explanation of the algorithm. To avoid tedious notation, we describe the two-site case but the idea behind the one-site additive algorithm is identical.
Essentially, the A2DMRG algorithm consists of the following steps: Given an initial low-rank tensor $\mathbb{Y}^{(0),0}\in\mathcal{T}_{\mathbb{r}}$, we first perform $d-1$ local optimization steps. For each $j\in\{1,\ldots,d-1\}$, the $j^{\rm th}$ local optimization step optimizes the combined $j^{\rm th}$ and $j+1^{\rm st}$ two-site TT core of the $j$-orthogonal TT decomposition of $\mathbb{Y}^{(0),0}$, this choice of orthogonalization being motivated by Lemma 10 and Remark 11. Consequently, each of these local steps can be performed independently in parallel and we obtain $d-1$ updated tensors $\{{\mathbb{Y}}^{(0),j}\}_{j=1}^{d-1}$. Next, we compute an optimal linear combination $\widetilde{\mathbb{Y}}^{(1),0}\in\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ of the tensors $\{{\mathbb{Y}}^{(0),j}\}_{j=0}^{d-1}$ by optimizing the given energy functional $\mathcal{J}\colon U\subset\mathbb{R}^{n_{1}\times\ldots\times n_{d}}\rightarrow\mathbb{R}$ on the coarse space $\text{span}\{{\mathbb{Y}}^{(0),j}\}_{j=0}^{d-1}$. Finally, we apply a compression step by projecting $\widetilde{\mathbb{Y}}^{(1),0}$ onto the low-rank manifold $\mathcal{T}_{\widetilde{\mathbb{r}}}\subset\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ for a judicious choice of rank parameters $\widetilde{\mathbb{r}}\in\mathbb{N}^{d-1}$. This yields an updated low-rank tensor $\mathbb{Y}^{(1),0}\in\mathcal{T}_{\widetilde{\mathbb{r}}}$ that can be used as a new initialization.
@@
Left-orthogonal tensor train decomposition $\mathbb{Y}_{d}^{(0)}\in\mathcal{V}_{\mathbb{r}}=\tau^{-1}\big{(}\mathcal{T}_{\mathbb{r}}\cap U\big{)}\subset{\mathcal{U}}_{\mathbb{r}}$.
@@
     Save a left-orthogonal TT decomposition $\mathbb{Y}_{d}^{(n)}=(\mathbb{U}_{1}^{(n)},\ldots,\mathbb{U}_{d-1}^{(n)},{\mathbb{W}}_{d}^{(n)})\in\mathcal{V}_{\mathbb{r}}$: {mdframed}[hidealllines=true,backgroundcolor=gray!20]
     for $i=d,d-1,\ldots,2$ do
         Given an $i$-orthogonal TT decomposition:
$\mathbb{Y}_{i}^{(n)}=\big{(}\mathbb{U}_{1}^{(n)},\ldots,\mathbb{U}_{i-1}^{(n)},\mathbb{W}_{i}^{(n)},\mathbb{V}_{i+1}^{(n)},\ldots,\mathbb{V}_{d}^{(n)}\big{)}\in\mathcal{V}_{\mathbb{r}}$
         Perform an orthogonalization step:
$\text{Compute }~{}{\mathbb{W}}_{i}^{<1>}=\mathbb{L}_{i}\mathbb{Q}_{i}\text{ with }\mathbb{L}_{i}\in\mathbb{R}^{r_{i-1}\times r_{i-1}},~{}\mathbb{Q}_{i}\in\mathbb{R}^{r_{i-1}\times n_{i}r_{i}},$
Save the $(i-1)$-orthogonal tensor train decomposition:
$\hskip 41.25641pt\mathbb{Y}_{i-1}^{(n)}=\big{(}\mathbb{U}_{1}^{(n)},\ldots,\mathbb{U}_{i-2}^{(n)},\mathbb{W}_{i-1}^{(n)},\mathbb{V}_{i}^{(n)},\ldots,\mathbb{V}_{d}^{(n)}\big{)}\in\mathcal{V}_{\mathbb{r}}$
$\hskip 41.25641pt\text{where }\mathbb{W}_{i-1}^{(n)}=\mathbb{U}_{i-1}^{(n)}\mathbb{L}_{i},~{}\mathbb{V}_{i}^{(n)}=\mathbb{Q}_{i}.$
     end for {mdframed}[hidealllines=true,backgroundcolor=gray!60]
         for $i=1,2,\ldots,d$ (respectively, $i=1,2,\ldots,d-1$) in parallel do
              Perform a one-site (respectively, two-site) DMRG micro-step, i.e.,
$\hskip 21.33955pt\text{Compute }~{}\mathbb{V}_{i}=\mathcal{S}_{i}(\mathbb{Y}_{i}^{(n)}),\text{ or compute }~{}\mathbb{W_{i,i+1}}=\mathcal{S}_{i,i+1}(\mathbb{Y}_{i}^{(n)}).$
              Define updated TT decomposition $\widetilde{\mathbb{Y}}_{i}^{(n+1)}$ using $\mathbb{V}_{i}$ or $\mathbb{W_{i,i+1}}$.
         end for {mdframed}[hidealllines=true,backgroundcolor=gray!20]
         Compute a minimizer $\{c^{*}_{j}\}_{j=0}^{d}\in\mathbb{R}^{d}$ of the minimization problem
(13) $\underset{\{c_{j}\}_{j=0}^{d}\in\mathbb{R}^{d}}{\text{min}}\;\mathcal{J}\Big{(}\sum_{j=1}^{d}c_{j}\tau\left(\widetilde{\mathbb{Y}}_{i}^{(n+1)}\right)\Big{)}.$
{mdframed}[hidealllines=true,backgroundcolor=gray!60]
     Apply compression: compute left-orthogonal, quasi-optimal approximation
$\mathcal{V}_{\widetilde{\mathbb{r}}}\ni\mathbb{Y}^{(n+1)}_{d}\approx\sum_{j=1}^{d}c^{*}_{j}\widetilde{\mathbb{Y}}_{j}^{(n+1)},\qquad\text{for a given choice of rank parameter }\widetilde{\mathbb{r}}\geq\mathbb{r}.$
@@
The first step of the A2DMRG Algorithm 2 consists a sequence of LQ orthogonalizations are performed on the components of the initial tensor train decomposition $\mathbb{Y}_{d}^{(n)}$. The goal of this step is to obtain a $j$-orthogonal copy of $\mathbb{Y}_{d}^{(n)}$ for each $j\in\{1,\ldots,d\}$. Note that this type of orthogonalization procedure is typically applied as a pre-processing step in the classical one-site and two-site DMRG algorithms 1 to obtain an appropriate initialization. In the A2DMRG Algorithm 2 therefore, we simply have to apply the orthogonalization twice. Note that current implementations of this step are sequential in nature.
In the second step, for each $j\in\{1,\ldots,d\}$ (respectively, $j\in\{1,\ldots,d-1\}$), we apply the $j^{\rm th}$ one-site (respectively, two-site) DMRG micro-iteration on the $j$-orthogonal copy of $\mathbb{Y}_{d}^{(n)}$. In view of Lemma 10, the resulting micro-iterations are numerically well-conditioned if the energy functional $\mathcal{J}$ is a quadratic form involving a well-conditioned symmetric positive definite tensor operator $\mathbb{A}$. In case the energy functional $\mathcal{J}$ is a Rayleigh quotient involving a symmetric operator $\mathbb{A}$, Remark 11 shows that the micro-iterations involve the solution of a symmetric eigenvalue problem. We emphasize that each one-site/two-site DMRG micro-iteration in this step can be performed by an independent processor in parallel, and this is discussed in further detail in Section A in the supplementary material.
In the third step, we perform a second-level minimization by seeking a minimizer of the energy functional $\mathcal{J}$ on a vector space of dimension at most $d$. This step, which is the subject of extensive discussion in the forthcoming Section 3.1 below, requires gathering the eigenfunctions (reshaped into updated TT cores) computed by each processor in parallel. Note that this is similar to two-level domain decomposition methods for solving linear systems of equations.
Since any minimizer obtained in the previous step will, in general, possess a tensor train decomposition with higher ranks than the initial ranks $\mathbb{r}$ of $\mathbb{Y}_{d}^{(n)}$, the fourth step in A2DMRG consists of the application of some compression algorithm (such as the TT rounding algorithm of Oseledets [33]) to this minimizer so as to obtain a quasi-optimal approximation with ranks $\mathbb{r}$. This step, which is also sequential in nature, is the subject of further discussion in Section A in the supplementary material.
@@
As noted in Lemma 10, for each $i\in\{1,\ldots,d\}$, the $i^{\rm th}$ one-site DMRG micro-iteration consists of using the Lanczos algorithm to compute the lowest eigenpair of a matrix $\mathbb{A}_{i}$ of dimension $r_{i-1}n_{i}r_{i}\times r_{i-1}n_{i}r_{i}$. The lowest eigenvector is then used to construct an updated TT decomposition by replacing the $i^{\rm th}$ TT core of the previous iterate with the newly obtained eigenvector (see Step 2 of Algorithm 2). The coarse space is thus constructed using these $d$ updated TT decompositions.
Notice, however, that the symmetric matrix $\mathbb{A}_{i}$ possesses $r_{i-1}n_{i}r_{i}$ orthogonal eigenvectors. An obvious idea is therefore to store not only the lowest eigenvector of $\mathbb{A}_{i}$ but the $K\in\mathbb{N}$ lowest eigenvectors yielded by the Lanczos algorithm. We can thus construct $K$ updated TT decompositions by replacing the $i^{\rm th}$ TT core of the previous iterate with the each of the newly obtained and reshaped $K$ eigenvectors. Proceeding in this manner for each of the $d$ one-site DMRG micro-iterations allows to construct an enriched coarse space using the linear span of $Kd$ tensors rather than $d$ tensors as before. Similar considerations, of course, apply to two-site A2DMRG.
@@
Given a TT decomposition consisting of $d$ TT cores, the one-site A2DMRG Algorithm 2 consists of performing $d$ independent, parallel local optimizations, one for each TT core. This yields $d$ updated TT decompositions and a coarse space consisting of the linear span of $d$ tensors.
@@
Another natural idea is to instead partition the $d$ TT cores into $M<d$ groups, with each group consisting of $d/M$ adjoining TT cores. For any such group, the $d/M$ adjoining TT cores can then be sequentially optimized using the classical DMRG algorithm, resulting in a single, new TT decomposition that contains $d/M$ updated TT cores. Proceeding in this manner for each of the $M$ groups yields therefore yields $M$ updated TT decompositions (each computed independently and in parallel) with a corresponding coarse space constructed using these $M$ updated TT decompositions. Note that this extension can, in particular, be seen as combining the approach of Stoudenmire and White [39] with a second-level coarse space correction.
A more detailed exploration of these ideas will be the subject of future work.
Remark 14 (Compression Step in Algorithm 2).
@@
A detailed study of the computational cost of the additive DMRG Algorithm 2 in comparison to the classical, sequential DMRG algorithm is the subject of Section A in the supplementary material. Let us nonetheless briefly comment on the computational cost of the final compression step of Algorithm 2.
@@
For the one-site (respectively, two-site) algorithm, this step consists of compressing the TT ranks of a linear combination of $d$ (respectively, $d-1$) tensors in TT format each with maximal rank ${r}\in\mathbb{N}$. It is well-known that such a linear combination of $d$ tensors possesses a tensor train decomposition of maximal rank $d{r}$. At first glance therefore, using the TT-rounding algorithm of Oseledets [33] or its various randomized improvements would then cost $\mathcal{O}(d^{3}nr^{3})$ flops but this is misleading.
@@
The essential point is that each of the TT decompositions in the linear combination differs from the rest in at most two (respectively, four) TT cores. Consequently, any tensor resulting from a linear combination of such TT format tensors is highly structured (for instance, in the one-site case, it has a TT decomposition of maximal rank $2r$). The compression step can therefore be performed with a much reduced cost. We refer to Section A in the supplementary material for details.

### v2

@@ 2.2. Alternating Minimization on the Tensor Train Parameter Space
$\hskip 28.45274pt\mathbb{U}_{\rm LR}^{i,(n+1)}=(\mathbb{U}^{(n+1)}_{1},\ldots,\mathbb{U}^{(n+1)}_{i-1},\widetilde{\mathbb{U}}_{i}^{(n)},\mathbb{U}^{(n)}_{i+1},\ldots,\mathbb{U}^{(n)}_{d})$
$\hskip 11.38092pt\hookrightarrow\mathbb{U}_{\rm LR}^{i+1,(n+1)}=(\mathbb{U}^{(n+1)}_{1},\ldots,\mathbb{U}^{(n+1)}_{i-1},\mathbb{U}^{(n+1)}_{i},\widetilde{\mathbb{U}}^{(n)}_{i+1},\ldots,\mathbb{U}^{(n)}_{d})$,
$\hskip 28.45274pt\text{where }\mathbb{U}^{(n+1)}_{i}=\mathbb{P}_{i},~\widetilde{\mathbb{U}}^{(n)}_{i+1}=\mathbb{S}_{i}\mathbb{Q}_{i}.$
@@ 3. Additive Two-Level DMRG with Coarse Space Correction
Equipped with the tools and notions presented in Section 2, we are now ready to present our proposal for the additive two-level DMRG algorithm with coarse space correction, or A2DMRG for short. Prior to presenting the details, let us first present a high-level intuitive explanation of the algorithm. To avoid tedious notation, we describe the two-site case but the idea behind the one-site version is identical.
Essentially, the A2DMRG algorithm consists of the following steps: Given an initial low-rank tensor $\mathbb{Y}^{(0),0}\in\mathcal{T}_{\mathbb{r}}$, we first perform d1d-1 local optimization steps. For each $j\in\{1,\ldots,d-1\}$, the $j^{\rm th}$ local optimization step optimizes the combined $j^{\rm th}$ and $j+1^{\rm st}$ two-site TT core of the $j$-orthogonal TT decomposition of $\mathbb{Y}^{(0),0}$, this choice of orthogonalization being motivated by Lemma 10 and Remark 11. Consequently, each of these local steps can be performed independently in parallel and we obtain d1d-1 updated tensors $\{{\mathbb{Y}}^{(0),j}\}_{j=1}^{d-1}$. Next, we compute an optimal linear combination $\widetilde{\mathbb{Y}}^{(1),0}\in\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ of the tensors $\{{\mathbb{Y}}^{(0),j}\}_{j=0}^{d-1}$ by optimizing the given energy functional $\mathcal{J}\colon U\subset\mathbb{R}^{n_{1}\times\ldots\times n_{d}}\rightarrow\mathbb{R}$ on the coarse space $\text{span}\{{\mathbb{Y}}^{(0),j}\}_{j=0}^{d-1}$. Finally, we apply a compression step by projecting $\widetilde{\mathbb{Y}}^{(1),0}$ onto the low-rank manifold $\mathcal{T}_{\widetilde{\mathbb{r}}}\subset\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ for a judicious choice of rank parameters $\widetilde{\mathbb{r}}\in\mathbb{N}^{d-1}$. This yields an updated low-rank tensor $\mathbb{U}^{(1),0}\in\mathcal{T}_{\widetilde{\mathbb{r}}}$ that can be used as a new initialization.
@@
Left-orthogonal tensor train decomposition $\mathbb{U}^{(0),d}\in\mathcal{V}_{\mathbb{r}}=\tau^{-1}\big(\mathcal{T}_{\mathbb{r}}\cap U\big)$.
@@
  Save a left-orthogonal TT decomposition $\mathbb{U}^{(n),d}=(\mathbb{U}_{1}^{(n)},\ldots,\mathbb{U}_{d-1}^{(n)},{\mathbb{W}}_{d}^{(n)})\in\mathcal{V}_{\mathbb{r}}$:   for $i=d,d-1,\ldots,2$ do    Given an $i$-orthogonal $T$ decomposition: $\mathbb{U}^{(n),i}=\big(\mathbb{U}_{1}^{(n)},\ldots,\mathbb{U}_{i-1}^{(n)},\mathbb{W}_{i}^{(n)},\mathbb{V}_{i+1}^{(n)},\ldots,\mathbb{V}_{d}^{(n)}\big)\in\mathcal{V}_{\mathbb{r}}$    Perform an orthogonalization step: $\text{Compute }~{\mathbb{W}}_{i}^{<1>}=\mathbb{L}_{i}\mathbb{Q}^{<i>}_{i}\text{ with }\mathbb{L}_{i}\in\mathbb{R}^{r_{i-1}\times r_{i-1}},~\mathbb{Q}^{<i>}_{i}\in\mathbb{R}^{r_{i-1}\times n_{i}r_{i}},$ Save the (i1)(i-1)-orthogonal tensor train decomposition: $\hskip 29.87547pt\mathbb{U}^{(n),{i-1}}=\big(\mathbb{U}_{1}^{(n)},\ldots,\mathbb{U}_{i-2}^{(n)},\mathbb{W}_{i-1}^{(n)},\mathbb{V}_{i}^{(n)},\ldots,\mathbb{V}_{d}^{(n)}\big)\in\mathcal{V}_{\mathbb{r}}$ $\hskip 41.25641pt\text{where }\mathbb{W}_{i-1}^{(n)}=\mathbb{U}_{i-1}^{(n)}\mathbb{L}_{i},~\mathbb{V}_{i}^{(n)}=\mathbb{Q}_{i}.$   end for    for $i=1,2,\ldots,d$ (respectively, $i=1,2,\ldots,d-1$) in parallel do      Perform a one-site (respectively, two-site) DMRG micro-step, i.e., $\hskip 32.72049pt\text{Compute }~\mathbb{V}_{i}=\mathcal{S}_{i}(\mathbb{U}^{(n),i}),\text{ or compute }~\mathbb{W_{i,i+1}}=\mathcal{S}_{i,i+1}(\mathbb{U}^{(n),i}).$      Define updated TT decomposition $\widetilde{\mathbb{U}}^{(n+1),i}$ using $\mathbb{V}_{i}$ or $\mathbb{W_{i,i+1}}$.    end for    Set $\widetilde{\mathbb{U}}^{(n+1),0}=\mathbb{U}^{(n),d}$.   Compute a minimizer $\{c^{*}_{j}\}_{j=0}^{d}\in\mathbb{R}^{d+1}$ of (13) $\underset{\{c_{j}\}_{j=0}^{d}\in\mathbb{R}^{d+1}}{\text{min}}\;\mathcal{J}\Big(\sum_{j=0}^{d}c_{j}\tau\left(\widetilde{\mathbb{U}}^{(n+1),j}\right)\Big)\qquad$ $\text{(One-Site Case)},$      or a minimizer $\{d^{*}_{j}\}_{j=0}^{d-1}\in\mathbb{R}^{d-1}$ of $\underset{\{d_{j}\}_{j=0}^{d-1}\in\mathbb{R}^{d}}{\text{min}}\;\mathcal{J}\Big(\sum_{j=0}^{d-1}d_{j}\tau\left(\widetilde{\mathbb{U}}^{(n+1),j}\right)\Big)\qquad$ $\text{(Two-Site Case)}.$   Apply compression, i.e., compute left-orthogonal, quasi-optimal approximation $\mathcal{V}_{\widetilde{\mathbb{r}}}\ni\mathbb{U}^{(n+1),d}\approx\sum_{j=0}^{d}c^{*}_{j}\widetilde{\mathbb{U}}^{(n+1),j}\qquad$ $\text{(One-Site Case)}\quad\text{or},$ $\mathcal{V}_{\widetilde{\mathbb{r}}}\ni\mathbb{U}^{(n+1),d}\approx\sum_{j=0}^{d-1}d^{*}_{j}\widetilde{\mathbb{U}}^{(n+1),j}\qquad$ $\text{(Two-Site Case)},$ for a given choice of rank parameter $\widetilde{\mathbb{r}}\geq\mathbb{r}$.
@@
The first step of the A2DMRG Algorithm 2 consists a sequence of LQ orthogonalizations are performed on the components of the initial tensor train decomposition $\mathbb{U}^{(n),d}$. The goal of this step is to obtain a $j$-orthogonal copy of $\mathbb{U}^{(n),d}$ for each $j\in\{1,\ldots,d\}$. Note that this type of orthogonalization procedure is typically applied as a pre-processing step in the classical one-site and two-site DMRG algorithms 1 to obtain an appropriate initialization. In the A2DMRG Algorithm 2 therefore, we simply have to apply the orthogonalization twice. Current implementations of this step are sequential in nature.
In the second step, for each $j\in\{1,\ldots,d\}$ (respectively, $j\in\{1,\ldots,d-1\}$), we apply the $j^{\rm th}$ one-site (respectively, two-site) DMRG micro-iteration on the $j$-orthogonal copy of $\mathbb{U}^{(n),d}$. In view of Lemma 10, the resulting micro-iterations are numerically well-conditioned if the energy functional $\mathcal{J}$ is a quadratic form involving a well-conditioned symmetric positive definite tensor operator $\mathbb{A}$. In case the energy functional $\mathcal{J}$ is a Rayleigh quotient involving a symmetric operator $\mathbb{A}$, Remark 11 shows that the micro-iterations involve the solution of a symmetric eigenvalue problem. We emphasize that each one-site/two-site DMRG micro-iteration in this step can be performed by an independent processor in parallel, and this is discussed in further detail in Section A in the supplementary material.
In the third step, we perform a second-level minimization by seeking a minimizer of the energy functional $\mathcal{J}$ on a vector space of dimension at most $d+1$ (one-site setting) or $d$ (two-site setting). This step, which is the subject of extensive discussion in Section 3.1 below, requires gathering the eigenfunctions (reshaped into updated TT cores) computed by each processor in parallel. Note that this is similar to two-level domain decomposition methods for solving linear systems of equations.
Since any minimizer obtained in the previous step will, in general, possess a tensor train decomposition with higher ranks than the initial ranks $\mathbb{r}$ of $\mathbb{U}^{(n),d}$, the fourth step in A2DMRG consists of the application of some compression algorithm (such as the TT rounding algorithm of Oseledets [35]) to this minimizer so as to obtain a quasi-optimal approximation with lower ranks. This step, which is– at least partially– also sequential in nature, is the subject of further discussion in Section A in the supplementary material.
@@
As noted in Lemma 10, for each $i\in\{1,\ldots,d\}$, the $i^{\rm th}$ one-site DMRG micro-iteration consists of using the Lanczos algorithm to compute the lowest eigenpair of a matrix $\mathbb{A}_{i}$ of dimension $r_{i-1}n_{i}r_{i}\times r_{i-1}n_{i}r_{i}$. An updated TT decomposition is then obtained by replacing the $i^{\rm th}$ TT core of the previous iterate with the newly computed lowest eigenvector (see Step 2 of Algorithm 2). The coarse space is constructed using these $d$ updated TT decompositions along with the previous global iterate.
Notice, however, that the symmetric matrix $\mathbb{A}_{i}$ possesses $r_{i-1}n_{i}r_{i}$ orthogonal eigenvectors. An obvious idea is therefore to store not only the lowest eigenvector of $\mathbb{A}_{i}$ but the $K\in\mathbb{N}$ lowest eigenvectors yielded by the Lanczos algorithm. We can then construct $K$ updated TT decompositions by replacing the $i^{\rm th}$ TT core of the previous iterate with the each of the newly obtained and reshaped $K$ eigenvectors. Proceeding in this manner for each of the $d$ one-site DMRG micro-iterations allows to construct an enriched coarse space using the linear span of $Kd+1$ tensors rather than $d+1$ tensors as before. Similar considerations, of course, apply to two-site A2DMRG.
@@
Given a TT decomposition consisting of $d$ TT cores, the one-site A2DMRG Algorithm 2 consists of performing $d$ independent, parallel local optimizations, one for each TT core. This yields $d$ updated TT decompositions and a coarse space consisting of the linear span of dd+1 tensors.
@@
Another natural idea is to instead partition the $d$ TT cores into $M<d$ groups, with each group consisting of $d/M$ adjoining TT cores. For any such group, the $d/M$ adjoining TT cores can then be sequentially optimized using the classical DMRG algorithm, resulting in a single, new TT decomposition that contains $d/M$ updated TT cores. Proceeding in this manner for each of the $M$ groups therefore yields $M$ updated TT decompositions (each computed independently and in parallel) with a corresponding coarse space constructed using these $M$ updated TT decompositions along with the previous global iterate. Note that this extension can, in particular, be seen as combining the approach of Stoudenmire and White [41] with a second-level coarse space correction.
An exploration of these ideas will be the subject of future work.
Remark 14 (Compression Step in A2DMRG Algorithm 2).
@@
A detailed study of the computational cost of the A2DMRG Algorithm 2 in comparison to the classical, sequential DMRG algorithm is the subject of Section A in the supplementary material. Let us nonetheless briefly comment on the computational cost of the final compression step of Algorithm 2.
@@
For the one-site (respectively, two-site) algorithm, this step consists of compressing the TT ranks of a linear combination of $d+1$ (respectively, $d$) tensors in TT format each with maximal rank ${r}\in\mathbb{N}$. It is well-known that such a linear combination of $d$ tensors possesses a tensor train decomposition of maximal rank (d+1)r(d+1){r} (respectively, $dr$). At first glance, using the TT-rounding algorithm of Oseledets [35] or its various randomized improvements would then cost $\mathcal{O}(d^{3}nr^{3})$ flops but this is misleading.
@@
The essential point is that each of the TT decompositions in the linear combination differs from the rest in at most two (respectively, four) TT cores. Consequently, any tensor resulting from a linear combination of such TT format tensors is highly structured (for instance, in the one-site case, it has a TT decomposition of maximal rank $2r$). The compression step can therefore be performed with a much reduced cost. We refer to Section A in the supplementary material for details. Table 1 below also summarizes the dominant computational cost of one global iteration (one half-sweep, respectively,) of the A2DMRG algorithm (classical DMRG algorithm, respectively).
Half-sweep DMRG cost A2MDRG global iteration cost per processor
1-site $\mathcal{O}\big(dnr^{3}R(K+d)\big)$ $\mathcal{O}\big(nr^{3}R(K+d)\big)+\mathcal{O}\big(dnr^{3}RK^{\prime}\big)$
2-site $\mathcal{O}\big(dnr^{3}R(nK+d)\big)$ $\mathcal{O}\big(nr^{3}R(Kn+d)\big)+\mathcal{O}\big(dn^{3}r^{3}RK^{\prime}\big)$
Table 1. Dominant computational cost of the classical DMRG and A2DMRG methods for solving eigenvalue problems arising from quantum chemistry. Here, $n=\max\{n_{j}\}_{j\in\mathbb{N}}$, $r=\max\mathbb{r}$, $R$ is the maximal rank in the MPO representation of the underlying tensor operator $\mathbb{A}$, $K$ is the maximal number of Krylov iterations required for all DMRG micro-iterations and $K^{\prime}$ is the number of Krylov iterations required for the solution of the second-level minimization problem. We also assume that $nR<r$, and that the SVD steps in two-site DMRG micro-iterations produce maximal ranks $\widetilde{r}=\mathcal{O}(r)$. For a derivation of these costs, we refer to the supplementary material.
Remark 15 (Connection to Additive Schwarz Methods).
As claimed in the introduction, the A2DMRG Algorithm 2 is inspired by the well-known additive Schwarz methods from the domain decomposition literature. To see this connection, let us briefly recall the abstract framework of additive Schwarz methods [15]: Let $\big(V,(\cdot,\cdot)_{V}\big)$ be a finite-dimensional Hilbert space, let $A\colon V\rightarrow V$ be a symmetric, positive definite operator, and suppose we are interested in solving, for some given $b\in V$
$Au$ $=b\qquad\text{or equivalently, in variational form,}$
$\underset{v\in V}{\text{argmin}}\mathcal{R}(v)$ $\qquad\text{with }~\mathcal{R}(v)=\frac{1}{2}(Av,v)_{V}-(b,v)_{V}.$
We assume that $V$ possesses a decomposition of the form $V=\sum_{j=1}^{d}R_{j}^{\dagger}V_{j}$ where each $V_{j}$ is a Hilbert space and $R_{j}^{\dagger}\colon V_{j}\rightarrow V$ is a prolongation operator with corresponding restriction operator $R_{j}\colon V\rightarrow V_{j}$. In this framework, assuming the initialization $u^{(0)}\in V$, an additive Schwarz method with damping factors $\{\omega_{j}\}_{j=0}^{d}\in\mathbb{R},~\omega_{0}=1,$ can be defined through the iterates
$u^{(\ell+1)}=\omega_{0}u^{(\ell)}-\sum_{j=1}^{d}\omega_{j}R_{j}^{\dagger}{z}^{(\ell)}_{j}\hskip 5.69054pt\text{with }~{z}^{(\ell)}_{j}=(R_{j}AR_{j}^{\dagger})^{-1}R_{j}Au^{(\ell)}-(R_{j}AR_{j}^{\dagger})^{-1}R_{j}b.$
The key observation is that each update $z^{(\ell)}_{j}\in V_{j}$ defined above can be viewed as the minimizer of an energy functional, i.e.,
$z^{(\ell)}_{j}$= $\underset{y_{j}\in V_{j}}{\text{argmin}}\frac{1}{2}\big(R_{j}AR_{j}^{\dagger}y_{j},y_{j}\big)_{V}-\big(R_{j}Au^{(\ell)}-R_{j}b,y_{j}\big)_{V}=\underset{y_{j}\in V_{j}}{\text{argmin}}\;\mathcal{R}(u^{(\ell)}-R_{j}^{\dagger}y_{j}),$
so that $~u^{(\ell+1)}=\omega_{0}u^{(\ell)}-\sum_{j=1}^{d}\omega_{j}R_{j}^{\dagger}\;\underset{y_{j}\in V_{j}}{\text{argmin}}\;\mathcal{R}(u^{(\ell)}-R_{j}^{\dagger}y_{j}).$
Using now a bit of algebra together with the multi-linearity of the tensor contraction mapping and known properties of the retraction operators [19, Lemma 3.2], we can deduce that one global iteration of the one-site A2DMRG algorithm 2 with initialization $\tau(\mathbb{U}^{(\ell)})\in\mathcal{T}_{\mathbb{r}}$ corresponds to a generalization of the above additive Schwarz iteration where the quadratic energy functional $\mathcal{R}$ is replaced with a general functional $\mathcal{J}\colon U\subset\mathbb{R}^{n_{1}\times\ldots n_{d}}\rightarrow\mathbb{R}$ and
$V$ $=\mathbb{T}_{\tau(\mathbb{U}^{(\ell)})}\mathcal{T}_{\mathbb{r}},\quad\text{i.e., the tangent space at $\tau(\mathbb{U}^{(\ell)})$ to $\mathcal{T}_{\mathbb{r}}$},$
$R_{j}^{\dagger}$ $=\mathbb{P}_{\mathbb{U}^{(\ell)},j,1},\qquad\text{and}\qquad V_{j}=\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}},$
$R_{j}^{\dagger}V_{j}$ $=\left\{\tau({\mathbb{U}}^{(\ell),j})\colon{\mathbb{U}}^{(\ell),j}=(\mathbb{U}^{(\ell)}_{1},\ldots,\mathbb{U}_{j-1}^{(\ell)},\mathbb{V}_{j},\mathbb{U}_{j+1}^{(\ell)},\ldots\mathbb{U}_{d}^{(\ell)}),~\mathbb{V}_{j}\in\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}\right\},$
$\{\omega_{j}\}_{j=0}^{d}\hskip 5.69054pt\text{ is determined by the second-level minimization}.$
A closer study of this connection as well as a convergence analysis of A2DMRG is work in progress.